# Tuning of altermagnetism by strain

### v1

@@ I Introduction
Altermagnetism is a novel type of magnetic order that shares some properties with ferromagnets and some with antiferomagnets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Similar to antiferromagnets, an altermagnet has zero nonrelativistic net magnetization ($\mathbf{M}$). At the same time, altermagnets exhibit a finite anomalous Hall effect normally observed in ferromagnets. This implies that the magnetic symmetry group does not include elements combining time reversal $\mathcal{T}$ with inversion $\mathcal{P}$ or with spatial translation $T_{\mathbf{t}}$ by some vector $\mathbf{t}$.
@@
Tuning of altermagnets by strain [4, 11, 12] can enable useful functionalities for spintronic applications, such as the ability to manipulate and switch altermagnetic domains via magnetostrictive or piezomagnetic Zeeman coupling. The structure of the third-rank linear piezomagnetic response tensor is determined by the magnetic point group of the material and fully tabulated [13]. Upon contraction with the Levi-Civita symbol, the same tensor structure describes strain-induced anomalous Hall conductivity [14]. However, magnetic point groups do not specifically distinguish altermagnets, which are defined by their nonrelativistic spin groups [15, 16]. It is, therefore, of interest to determine what can be learned about strain-induced magnetization in altermagnets and its mechanisms from their spin group structure and to connect such symmetry considerations with specific materials via first-principles calculations.
@@ II.2 Band filling effect: Metallic Lieb lattice
We will illustrate the effect using a generalization of the minimal model of Ref.  27. At a finite strain $\varepsilon/2=\varepsilon_{xx}=-\varepsilon_{yy}$, the original square lattice with the basis $\mathbf{a}_{x}=(1,0)$, $\mathbf{a}_{y}=(0,1)$ is deformed into a rectangular lattice, with the basis, $\mathbf{a}^{\prime}_{x}=(1+\varepsilon/2,0)/\sqrt{1-(\varepsilon/2)^{2}}$, $\mathbf{a}^{\prime}_{y}=(0,1-\varepsilon/2)/\sqrt{1-(\varepsilon/2)^{2}}$. We formulate the effective band Hamiltonian, $H_{b}$ as a function of the components, $k^{\prime}_{x,y}$ of a Bloch vector $\mathbf{k}$ in the basis modified by strain, $\mathbf{b}^{\prime}_{x}=(1-\varepsilon/2,0)/\sqrt{1-(\varepsilon/2)^{2}}$, $\mathbf{b}^{\prime}_{y}=(0,1+\varepsilon/2)/\sqrt{1-(\varepsilon/2)^{2}}$ formally restoring the Brillouin zone to a square. As a result, we obtained an extension of the model of Ref. [27] in the presence of strain:
@@
This expression holds at $E_{F}\ll t_{2}$ and far from the Lifshitz transition, $t_{2d}\ll t_{2}$. Under these conditions, the Fermi contours look as shown in Fig. 2a, and the induced magnetization is dominated by the strain-induced band splitting, $M\approx\Delta E_{\Gamma}\,\hat{\mathbf{L}}/(2\pi t_{2})$, similarly to the two-valley model considered in Ref. 4. In general, however, the second term in Eq. (5) includes an additional contribution arising from the strain-induced effective mass anisotropy proportional to $\delta t_{2d}$, which manifests itself as shown in Fig. 2b.
The nonrelativistic band-filling mechanism of piezomagnetism described in this section is inactive in insulating altermagnets. However, strain-induced splittings of the Kramers doublets at the $\Gamma$ point are described by the same invariants, listed in Table 1, as the magnetization in the metallic case. For example, in rutile MnF2, which has the same ${}^{2}4/{}^{1}m{}^{2}m_{d}{}^{1}m_{y}$ SLG as the Lieb lattice, the $\varepsilon_{xy}$ strain linearly splits the Kramers doublets at $\Gamma$, which can be observed spectroscopically. Figure 3 illustrates this splitting by a first-priniples calculation for MnF2 for the valence band lying about 0.4 eV below the valence band maximum.
@@ II.3 Temperature-dependent contribution: Insulating transition-metal fluorides
The agreement with the exchange parameters obtained from fitting the experimental magnon spectra is good for MnF2 and FeF2. Reference 54 reported considerably larger $J_{2}$ parameters, which can be attributed to their use of a smaller $U=3$ eV. For CoF2, the exchange parameters agree with an earlier calculation [55] but the large value of $J_{1}$ disagrees with experiment. We attribute this discrepancy to the deficiency of the single-reference DFT$+U$ description of the partially filled $t_{2g}$ subshell of the Co2+ ion. We also noted that the relatively small $J_{1}$ in MnF2 changes sign if the PBE [56] functional is used instead of PBEsol [57], suggesting the DFT prediction for this parameter is unreliable. However, this difference has almost no effect on the piezomagnetic response.
@@ III.2 Strain-induced Dzyaloshinskii-Moriya interaction
Figure 7 shows the calculated total energies $E(M_{c})$ for a small applied strain $\varepsilon_{yz}=\varepsilon/2$ (for the fluorides) or $\varepsilon_{xx}=\varepsilon$ (for MnTe and CrSb) with $\varepsilon=0.02$. In MnF2 and MnTe the spin canting is below the sensitivity of our calculations, which may be explained by the suppressed orbital moment of the Mn2+ ions with the half-filled $3d_{\uparrow}^{5}$ configuration. This is in agreement with a very small experimental value for MnF2. A tiny $\Lambda_{14}$ of less than 0.01 $\mu_{B}$/f.u. can also be inferred from the calculation of Ref. 66. In contrast, strain-induced magnetization is substantial in FeF2, CoF2, and CrSb.
@@ IV Strain-induced non-unitary triplet superconductivity
Recent works on manifestations of the altermagnetism in superconductors include the $\pi$ transition in Josephson junction with the altermagnetic normal region [67, 68]. Superconducting altermagnets are argued to exhibit a specific non-reciprocal critical current, also known as superconducting diode effect, if parity is broken, for example, by an external electric field [69].
Depending on the parameters, the superconducting order parameter favored by the altermagnet with finite spin-orbit interaction can be singlet, triplet or a mixture of the two that breaks $\mathcal{T}$ [21]. Here we address the non-relativistic limit, where the emergent inversion symmetry inherent in SLGs allows to disentangle singlet and triplet Cooper pair correlations. As the $\mathcal{T}$ symmetry is broken by the magnetism such symmetry protects the equal spin Cooper pairs, [17]. These triplet pairs fall into two varieties distinguished by the total angular momentum, $m_{s}=\pm 1$ along $\mathbf{L}$. The nontrivial spin symmetry $[C_{2\hat{\mathbf{L}}_{\perp}}||\mathcal{A}]$ transforms one type of triplet pairs to another. This ensures that such superconducting state is unitary on average.
@@
The irreducible representations (irreps) of the spin group are obtained in App. A. They can be derived from the irreps of the site group, $A_{g(u)}$, $B_{1g(u)}$, $B_{2g(u)}$ and $B_{3g(u)}$, where parity even (odd) irreps are labeled by the $g$ ($u$) subscript. We specify the irreps by stating the representation of a non-trivial element $[C_{2\hat{\mathbf{L}}_{\perp}}||\mathcal{A}]$, and a generic element, $g=[C_{\phi\hat{\mathbf{L}}}||h]\in[\mathbf{C}_{\infty}||\mathbf{H}]$, where $h\in\mathbf{H}$, and $C_{\phi\hat{\mathbf{L}}}$ is a rotation around $\mathbf{L}$ by angle $\phi$.
@@
An example of the fermion operator transforming trivially, i.e. as $\mathbf{\Gamma}^{0;+}_{A_{g}}$ is the Hamiltonian. The scalar product $\mathbf{L}\cdot\mathbf{M}$, as well as the functions $f_{\mathrm{P}}(\boldsymbol{\varepsilon})$ in the Tab. 1 change sign under $[C_{2\hat{\mathbf{L}}_{\perp}}||\mathcal{A}]$, and therefore, transform as $\mathbf{\Gamma}^{0;-}_{A_{g}}$. Since the spin rotations $\mathbf{C}_{\infty}$ are represented trivially by Eq. (11), the addition of $[C_{2\hat{\mathbf{L}}_{\perp}}||\mathcal{A}]$ extends the site group $\mathbf{H}$ to the crystallographic point group, $D_{4h}$, $\mathbf{\Gamma}^{0;+}_{A_{g}}$ is $A_{1g}$ and $\mathbf{\Gamma}^{0;-}_{A_{g}}$ is $B_{2g}$ irreps of $D_{4h}$ [23].
@@
where $m^{\prime}_{s}\in\mathbb{N}$, and $m_{s}\in\mathbb{Z}$.
We now apply these result to classify the Cooper pair operators. Generally, the equal-spin triplets, $P^{t}_{\pm 1}$ may form four two-dimensional irreps, $\mathbf{\Gamma}^{\pm 1}_{B_{2u}}$, $\mathbf{\Gamma}^{1}_{A_{u}}$ and $\mathbf{\Gamma}^{1}_{B_{1u}}$, Eq. (IV). If the spin wave-functions of a Cooper pairs are $P^{t}_{0}$ or $P^{s}$, they form one of the four one-dimensional irreps $\mathbf{\Gamma}^{0;\pm}_{A_{g(u)}}$ ($\mathbf{\Gamma}^{0;\pm}_{B_{1g(u)}}$) when the corresponding $\varphi(\mathbf{k})\in A_{g(u)}(B_{1g(u)})$, respectively, Eq. (11). Otherwise, they give rise to the two-dimensional irrep $\mathbf{\Gamma}^{0}_{B_{2g(u)}}$.
@@
IV.0.1 Cooper pairs on the Lieb lattice
@@
IV.0.2 Cooper pairs in the rutile structure
@@
IV.1 Effect of strain on equal-spin triplet superconductivity
@@
Previously we have classified the magnetization along Néel vector, $M_{L}=\mathbf{M}\cdot\mathbf{L}$ as well as the strain components listed in Tab. 1 as $\mathbf{\Gamma}^{0,-}_{A_{g}}$, see Eq. (11). Based to Eq. (17) they couple to the $\mathbf{\Gamma}^{0,-}_{A_{g}}$ component listed in Eq. (IV.1). Accidentally this argument also shows that they couple to each other in agreement with the results of Sec. II.
@@
The first term of Eq. (IV.1) shows that the strain splits the $T_{c0}^{B_{1u}}$ by an amount $\delta T_{c}$ proportional to the strain. This change originate from the change of the density of states for a given band of a bandwidth, $W$. It can be estimated as $\Delta T_{c}/T_{c0}^{B_{1u}}\propto\Delta E_{\Gamma}/W$, where the strain induced splitting at $\Delta E_{\Gamma}$ has been introduced in Sec. II.2.
The spin state wave-function of the Cooper pairs is encoded in the $\mathbf{d}$-vector, $(\mathbf{d}_{\mathbf{k}}\cdot\boldsymbol{\sigma})i\sigma_{y}$ [71]. According to Eq. (IV.0.2) the two types of the Cooper pairs transforming as $\mathbf{\Gamma}^{1}_{B_{1u}}$ can be assigned a sublattice index. The two $\mathbf{d}$-vectors read $\mathbf{d}_{\mathbf{k},A}=\phi^{B_{1u}}_{\mathbf{k}}(\hat{X}+i\hat{Y})$ and $\mathbf{d}_{\mathbf{k},B}=\phi^{B_{1u}}_{C_{2x}^{-1}\mathbf{k}}(\hat{X}-i\hat{Y})$ [22]. The two sets of Cooper pairs carry the opposite spin angular momentum, $\mathbf{S}_{\mathbf{k},A}=i\mathbf{d}_{\mathbf{k},A}\times[\mathbf{d}_{\mathbf{k},A}]^{*}=|\phi^{B_{1u}}_{\mathbf{k}}|^{2}\hat{Z}$ and similarly, $\mathbf{S}_{\mathbf{k},b}=-|\phi^{B_{1u}}_{C_{2x}^{-1}\mathbf{k}}|^{2}\hat{Z}$. We note that unlike in the conventional relativistic classification, $\mathbf{d}_{\mathbf{k}}$ is detached from the real space transformations and does not transform as a vector. Instead, the pairs can be thought of as colored differently for the two-component order parameters or the by the same color for a single-component ones.
At zero strain and magnetization, $|\Delta_{1}^{B_{1u}}|=|\Delta_{2}^{B_{1u}}|$, and the total angular momentum carried by the electrons of the two partners cancel each other out. The superconductivity hence is on average unitary. It follows from Eq. (IV.1), that certain strain disrupts this balance, $|\Delta_{1}^{B_{1u}}|\neq|\Delta_{2}^{B_{1u}}|$ and turns the superconducting state globally non-unitary.
The second term in the free energy, Eq. (IV.1) shows that the magnetization $\mathbf{M}_{\perp}\perp\mathbf{L}$ imposes a two-fold variation of the relative phase of the two partner order parameters as $\mathbf{M}_{\perp}$ rotates. Such a magnetization can be induced by a weak field polarizing the itinerant carriers. Potentially, this may play a role in phase-sensitive superconducting systems.

### v2

@@ I Introduction
Altermagnetism is a novel type of magnetic order that shares some properties with ferromagnets and some with antiferromagnets (see reviews [1, 2, 3, 4] and references therein). Similar to antiferromagnets, an altermagnet has zero net magnetization ($\mathbf{M}$) in the nonrelativistic limit. At the same time, altermagnets exhibit a finite anomalous Hall effect normally observed in ferromagnets. This implies that the magnetic symmetry group does not include elements combining time reversal $\mathcal{T}$ with inversion $\mathcal{P}$ or with spatial translation $T_{\mathbf{t}}$ by some vector $\mathbf{t}$.
@@
Tuning of altermagnets by strain [5, 6, 7, 8, 9, 10] can enable useful functionalities for spintronic applications, such as the ability to manipulate and switch altermagnetic domains via magnetostrictive or piezomagnetic Zeeman coupling. The structure of the third-rank linear piezomagnetic response tensor is determined by the magnetic point group of the material and fully tabulated [11]. Upon contraction with the Levi-Civita symbol, the same tensor structure describes strain-induced anomalous Hall conductivity [9]. However, magnetic point groups do not specifically distinguish altermagnets, which are defined [12] by their nonrelativistic spin groups [13, 14]. It is, therefore, of interest to determine what can be learned about strain-induced magnetization in altermagnets and its mechanisms from their spin group structure and to connect such symmetry considerations with specific materials via first-principles calculations.
@@ II.2 Band filling effect: Metallic Lieb lattice
We will illustrate the effect using a generalization of the minimal model of Ref.  25. At a finite strain $\varepsilon/2=\varepsilon_{xx}=-\varepsilon_{yy}$, the original square lattice with the basis $\mathbf{a}_{x}=(1,0)$, $\mathbf{a}_{y}=(0,1)$ is deformed into a rectangular lattice, with the basis, $\mathbf{a}^{\prime}_{x}=(1+\varepsilon/2,0)/\sqrt{1-(\varepsilon/2)^{2}}$, $\mathbf{a}^{\prime}_{y}=(0,1-\varepsilon/2)/\sqrt{1-(\varepsilon/2)^{2}}$. We formulate the effective band Hamiltonian, $H_{b}$ as a function of the components, $k^{\prime}_{x,y}$ of a Bloch vector $\mathbf{k}$ in the basis modified by strain, $\mathbf{b}^{\prime}_{x}=(1-\varepsilon/2,0)/\sqrt{1-(\varepsilon/2)^{2}}$, $\mathbf{b}^{\prime}_{y}=(0,1+\varepsilon/2)/\sqrt{1-(\varepsilon/2)^{2}}$ formally restoring the Brillouin zone to a square. As a result, we obtained an extension of the model of Ref. [25] in the presence of strain:
@@
This expression holds at $E_{F}\ll t_{2}$ and far from the Lifshitz transition, $t_{2d}\ll t_{2}$. Under these conditions, the Fermi contours look as shown in Fig. 2a, and the induced magnetization is dominated by the strain-induced band splitting, $M\approx\Delta E_{\Gamma}\,\hat{\mathbf{L}}/(2\pi t_{2})$, similarly to the two-valley model considered in Ref. 5. In general, however, the second term in Eq. (5) includes an additional contribution arising from the strain-induced effective mass anisotropy proportional to $\delta t_{2d}$, which manifests itself as shown in Fig. 2b.
The nonrelativistic band-filling mechanism of piezomagnetism described in this section is inactive in insulating altermagnets. However, strain-induced splittings of the Kramers doublets at the $\Gamma$ point are described by the same invariants, listed in Table 1, as the magnetization in the metallic case. For example, in rutile MnF2, which has the same ${}^{2}4/{}^{1}m{}^{2}m_{d}{}^{1}m_{y}$ SLG as the Lieb lattice, the $\varepsilon_{xy}$ strain linearly splits the Kramers doublets at $\Gamma$, which can be observed spectroscopically. Figure 3 illustrates this splitting by a first-principles calculation for MnF2 for the valence band lying about 0.4 eV below the valence band maximum. Essentially, finite $\varepsilon_{xy}$ strain turns MnF2 into a Luttinger-compensated [2] ferrimagnetic insulator where spin splitting at the $\Gamma$ point is always allowed [44].
@@ II.3 Temperature-dependent contribution: Insulating transition-metal fluorides
The agreement with the exchange parameters obtained from fitting the experimental magnon spectra is good for MnF2 and FeF2. Reference 53 reported considerably larger $J_{2}$ parameters, which can be attributed to their use of a smaller $U=3$ eV. For CoF2, the exchange parameters agree with an earlier calculation [54] but the large value of $J_{1}$ disagrees with experiment. We attribute this discrepancy to the deficiency of the single-reference DFT$+U$ description of the partially filled $t_{2g}$ subshell of the Co2+ ion. We also noted that the relatively small $J_{1}$ in MnF2 changes sign if the PBE [55] functional is used instead of PBEsol [56], suggesting the DFT prediction for this parameter is unreliable. However, this difference has almost no effect on the piezomagnetic response.
@@ III.2 Strain-induced Dzyaloshinskii-Moriya interaction
Figure 7 shows the calculated total energies $E(M_{c})$ for a small applied strain $\varepsilon_{yz}=\varepsilon/2$ (for the fluorides) or $\varepsilon_{xx}=\varepsilon$ (for MnTe and CrSb) with $\varepsilon=0.02$. In MnF2 and MnTe the spin canting is below the sensitivity of our calculations, which may be explained by the suppressed orbital moment of the Mn2+ ions with the half-filled $3d_{\uparrow}^{5}$ configuration. This is in agreement with a very small experimental value for MnF2. A tiny $\Lambda_{14}$ of less than 0.01 $\mu_{B}$/f.u. can also be inferred from the calculation of Ref. 65. In contrast, strain-induced magnetization is substantial in FeF2, CoF2, and CrSb.
@@ IV Strain-induced non-unitary triplet superconductivity
In several recent works various manifestations of the altermagnetism in superconductors such as the current-phase relation of a Josephson junction with the altermagnetic normal region have been explored, [66, 67, 68, 69, 70]. Superconducting altermagnets are argued to exhibit a specific non-reciprocal critical current, also known as superconducting diode effect, if parity is broken, for example, by an external electric field [71].
Depending on the parameters, the superconducting order parameter favored by the altermagnet with finite spin-orbit interaction can be singlet, triplet or a mixture of the two that breaks $\mathcal{T}$ [19]. Here we address the non-relativistic limit, where the emergent inversion symmetry inherent in SLGs allows to disentangle singlet and triplet Cooper pair correlations [72, 73]. As the $\mathcal{T}$ symmetry is broken by the magnetism such symmetry protects the equal spin Cooper pairs, [15]. These triplet pairs fall into two varieties distinguished by the total angular momentum, $m_{s}=\pm 1$ along $\mathbf{L}$. The nontrivial spin symmetry $[C_{2\hat{\mathbf{L}}_{\perp}}||\mathcal{A}]$ transforms one type of triplet pairs to another. This ensures that such superconducting state is unitary on average.
@@
Recently, the irreducible representations (irreps) of spin point groups have been tabulated [75]. It was stressed that in cases where the spin-only group is continuous, such as in collinear altermagnets, an irrep carries the index of the site group irrep and an integer or half-integer label $m_{s}$ of the irrep of the spin-only group. As a result, when the spin-only group is represented by nonzero phases the irreps of the full spin group are distinct from the irreps of the parent point group, and, in general, a priori unknown. We argue that such is the case of the equal-spin triplet Cooper pairs carrying an integer angular momentum, $m_{s}=\pm 1$. Hence, we are naturally led to the need to find the SLG irreps carrying integer angular momentum.
In App. A we implement a modified procedure to identify the irreps of the ${}^{2}4/{}^{1}m{}^{2}m_{y}{}^{1}m_{d}$ spin group, which allows us to obtain the representation matrices along with the symmetry-adapted partner functions. In the present context, the latter are interpreted as the Cooper pair wave function.
The irreps carry the index of the site group: $A_{g(u)}$, $B_{1g(u)}$, $B_{2g(u)}$ or $B_{3g(u)}$, where parity-even and odd irreps are labeled by the subscript $g$ or $u$. We specify the irreps by stating the representation of a nontrivial element $[C_{2\hat{\mathbf{L}}_{\perp}}||\mathcal{A}]$, and a generic element $g=[C_{\phi\hat{\mathbf{L}}}||h]\in[\mathbf{C}_{\infty}||\mathbf{H}]$, where $h\in\mathbf{H}$ and $C_{\phi\hat{\mathbf{L}}}$ is a rotation around $\mathbf{L}$ by angle $\phi$.
@@
An example of the fermion operator transforming trivially, i.e. as $\mathbf{\Gamma}^{0;+}_{A_{g}}$ is the Hamiltonian. The scalar product $\mathbf{L}\cdot\mathbf{M}$, as well as the functions $f_{\mathrm{P}}(\boldsymbol{\varepsilon})$ in the Tab. 1 change sign under $[C_{2\hat{\mathbf{L}}_{\perp}}||\mathcal{A}]$, and therefore, transform as $\mathbf{\Gamma}^{0;-}_{A_{g}}$. Since in Eq. (11) the spin rotations $\mathbf{C}_{\infty}$ are represented just by a number 1, it defines two irrep of the non-trivial part of the spin group, which is in turn isomorphic to the parent crystallographic point group, $D_{4h}$. In this way, e.g. $\mathbf{\Gamma}^{0;+}_{A_{g}}$ is $A_{1g}$ and $\mathbf{\Gamma}^{0;-}_{A_{g}}$ is $B_{2g}$ irreps of $D_{4h}$ [21].
@@
where $m^{\prime}_{s}\in\mathbb{N}$, and $m_{s}\in\mathbb{Z}$. Both Eqs. (11) and (IV) are in agreement with Ref. [75]. For instance, the matrices of the irrep $\mathbf{\Gamma}^{m_{s}}_{B_{2u}}$ is related to the matrices, $D^{(\mu,5)}$ with $\mu=m_{s}$ by a simple similarity transformation exchanging the two representation partners.
Finally, as found in Ref. [75] by applying the Dimmock test, the anti-unitary operations do not lead to doubling of the representation space. Physically, this is because the anti-unitary operation $\mathcal{T}[C_{2\hat{\mathbf{L}}_{\perp}}||E]$ leaves the spin orientation unchanged. It follows that under the $\mathcal{T}[C_{2\hat{\mathbf{L}}_{\perp}}||E]$ operation the partner wave functions for spin singlets and triplets acquire at most a phase factor. Hence, no doubling of the dimension occurs in the present example.
This conclusion agrees with Ref. [23] finding no spectrum degeneracy doubling due to the anti-unitary symmetry in the case of a $d$-wave altermagnet candidate FeSb2 with the ${}^{2}m_{z}{}^{2}m_{y}{}^{1}m_{x}$ SLG. This has been confirmed by using the Wigner criterion, which is analogous to the Dimmock test, to the little group at various Bloch momenta.
We are now ready to apply these result to classify the Cooper pair wave functions. Generally, the equal-spin triplets, $P^{t}_{\pm 1}$ may form four two-dimensional irreps, $\mathbf{\Gamma}^{\pm 1}_{B_{2u}}$, $\mathbf{\Gamma}^{1}_{A_{u}}$ and $\mathbf{\Gamma}^{1}_{B_{1u}}$, Eq. (IV). If the spin wave-functions of a Cooper pairs are $P^{t}_{0}$ or $P^{s}$, they form one of the four one-dimensional irreps $\mathbf{\Gamma}^{0;\pm}_{A_{g(u)}}$ ($\mathbf{\Gamma}^{0;\pm}_{B_{1g(u)}}$) when the corresponding $\varphi(\mathbf{k})\in A_{g(u)}(B_{1g(u)})$, respectively, Eq. (11). Otherwise, they give rise to the two-dimensional irrep $\mathbf{\Gamma}^{0}_{B_{2g(u)}}$.
@@
IV.1 Cooper pairs on the Lieb lattice
@@
IV.2 Cooper pairs in the rutile structure
@@
IV.3 Effect of strain on equal-spin triplet superconductivity
@@
Previously we have classified the magnetization along Néel vector, $M_{L}=\mathbf{M}\cdot\mathbf{L}$ as well as the strain components listed in Tab. 1 as $\mathbf{\Gamma}^{0,-}_{A_{g}}$, see Eq. (11). Based to Eq. (17) they couple to the $\mathbf{\Gamma}^{0,-}_{A_{g}}$ component listed in Eq. (IV.3). Accidentally this argument also shows that they couple to each other in agreement with the results of Sec. II.
@@
The first term of Eq. (IV.3) shows that the strain splits the $T_{c0}^{B_{1u}}$ by an amount $\delta T_{c}$ proportional to the strain. This change originate from the change of the density of states for a given band of a bandwidth, $W$. It can be estimated as $\Delta T_{c}/T_{c0}^{B_{1u}}\propto\Delta E_{\Gamma}/W$, where the strain induced splitting at $\Delta E_{\Gamma}$ has been introduced in Sec. II.2.
The spin state wave-function of the Cooper pairs is encoded in the $\mathbf{d}$-vector, $(\mathbf{d}_{\mathbf{k}}\cdot\boldsymbol{\sigma})i\sigma_{y}$ [76]. According to Eq. (IV.2) the two types of the Cooper pairs transforming as $\mathbf{\Gamma}^{1}_{B_{1u}}$ can be assigned a sublattice index. The two $\mathbf{d}$-vectors read $\mathbf{d}_{\mathbf{k},A}=\phi^{B_{1u}}_{\mathbf{k}}(\hat{X}+i\hat{Y})$ and $\mathbf{d}_{\mathbf{k},B}=\phi^{B_{1u}}_{C_{2x}^{-1}\mathbf{k}}(\hat{X}-i\hat{Y})$ [20]. The two sets of Cooper pairs carry the opposite spin angular momentum, $\mathbf{S}_{\mathbf{k},A}=i\mathbf{d}_{\mathbf{k},A}\times[\mathbf{d}_{\mathbf{k},A}]^{*}=|\phi^{B_{1u}}_{\mathbf{k}}|^{2}\hat{Z}$ and similarly, $\mathbf{S}_{\mathbf{k},b}=-|\phi^{B_{1u}}_{C_{2x}^{-1}\mathbf{k}}|^{2}\hat{Z}$. We note that unlike in the conventional relativistic classification, $\mathbf{d}_{\mathbf{k}}$ is detached from the real space transformations and does not transform as a vector. Instead, the pairs can be thought of as colored differently for the two-component order parameters or the by the same color for a single-component ones.
At zero strain and magnetization, $|\Delta_{1}^{B_{1u}}|=|\Delta_{2}^{B_{1u}}|$, and the total angular momentum carried by the electrons of the two partners cancel each other out. The superconductivity hence is on average unitary. It follows from Eq. (IV.3), that certain strain disrupts this balance, $|\Delta_{1}^{B_{1u}}|\neq|\Delta_{2}^{B_{1u}}|$ and turns the superconducting state globally non-unitary.
The second term in the free energy, Eq. (IV.3) shows that the magnetization $\mathbf{M}_{\perp}\perp\mathbf{L}$ imposes a two-fold variation of the relative phase of the two partner order parameters as $\mathbf{M}_{\perp}$ rotates. Such a magnetization can be induced by a weak field polarizing the itinerant carriers. Potentially, this may play a role in phase-sensitive superconducting systems.